burp4801:

Circle A has a diameter of 7 inches, a circumference of 21.98 inches, and an area of 38.465 square inches. The diameter of circle B is 6 inches, the circumference is 18.84 inches, and the area is 28.26 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

burp4801:

kinda stuck atm

YRJ8498:

Welcome to Question Cove! What part are you stuck on?

burp4801:

thank you..im kinda stuck with part a and b I just blanked out on how to solve these problems

Questionnare:

I did this just last thursday and I got a 50% on it so idk if you want my help

burp4801:

oh i don't mind...I just want like some tips you know?

Questionnare:

well my teacher did give me feedback so I may be able to help

burp4801:

oh that would be good ngl

Questionnare:

So, for part A, you have to solve for pi.

burp4801:

using the 3.14 thingy?

YRJ8498:

The formula for the circumference of circle when Diameter is given = πD π = Circumference / Diameter For Circle A

YRJ8498:

A has a diameter of 7 inches, a circumference of 21.98 inches. π = 21.98 inches/7 inches π = 3.14

burp4801:

oh wait its diameter not radius , i see so the equation would be c= 2 pi*d ?

burp4801:

we need to solve the value of pi for each circle...thats what confuses me

surjithayer:

diameter=2r=7 cm \[circumference=2 \pi r=21.98\] \[2r \pi=21.98\] \[\pi=21.98/7=?\]

surjithayer:

r=7/2 \[\pi (\frac{ 7 }{ 2 })^2=38.465\] \[\pi=38.465\times (\frac{ 4 }{ 49 })=3.14 \] similarly solve for B.